# On the Constant Depth Implementation of Pauli Exponentials

### v3

@@ Abstract
We decompose for the first time, under the very restrictive linear nearest-neighbour connectivity, $Z\otimes Z\ldots\otimes Z$ exponentials of arbitrary length into circuits of constant depth using $\mathcal{O}(n)$ ancillae and two-body XX and ZZ interactions. Consequently, a similar method works for arbitrary Pauli exponentials. We prove the correctness of our approach, after introducing novel rewrite rules for circuits which benefit from qubit recycling. The decomposition has a wide variety of applications ranging from the efficient implementation of fault-tolerant lattice surgery computations, to expressing arbitrary stabilizer circuits via two-body interactions only, and to reducing the depth of NISQ computations, such as VQE.
Introduction: —
@@
Without being generalized to arbitrary rotation angles or number of qubits, constructions similar to ours appeared in [2, 3]. Those constructions were derived and verified by applying the ZX-calculus on small scale diagrams. In contrast, we introduce novel circuit rewrite rules, derive and show the correctness of our decomposition using circuit diagrams. Circuit extraction from ZX-diagrams is very challenging and known to be a hard problem [18]. Therefore, our approach of generalising Pauli exponential decompositions directly via circuit rewrite rules is scalable and advantageous.
@@
Measurement-based implementation: —
When implementing the Pauli exponentials on fault-tolerant machines, the $\theta$ rotation is first decomposed into a sequence of rotation gates supported by the QECC. This increases the depth proportionally to the length of the gate sequence. Moreover, the depth will additionally increase by a constant factor, as each rotation is probabilistic and might require corrections [4] (e.g. if the X-measurement results in the -1 eigenvalue, then a corrective rotation of $2\theta$ is needed). Corrections can be tracked in a Pauli frame, when the angles correspond to Pauli gates.
@@
Decomposing the Pauli exponentials: —
@@
The general compilation strategy for lattice surgery is to first decompose arbitrary algorithms into generalized Pauli exponentials [4], and then to implement these exponentials via multi-body measurements. This approach holds multiple disadvantages.
Firstly, the resulting Pauli exponentials become very long range and cover large portions of the layout. This hinders the parallelisation of the operations, and alternative compilation methods have been proposed to overcome this, such as [16] (shows how to perform a constant depth CNOT using two-body measurements), and [29], which compiles to Clifford+T instead of exponentiated Paulis.
@@
Secondly, long range multi-body interactions might introduce high weight correlated logical errors into the computation. Such correlated errors are difficult to decode. The advantage of using pairwise interactions is in the decoding of logical computations [17]. Nearest neighbour pairwise interactions will localize the correlated errors and will limit their weight to at most two.
Constant Depth Fault-Tolerant Protocols: —
Recently, there has been an increased interest in using only two-body interactions to implement error-correction (e.g. [3, 2]). Two-body interactions are preferred as quantum computers have limited connectivity between the qubits, and very often the connectivity is not long range (e.g. [19]).
From the perspective of stabilizer QECC codes, our decomposition allows for any set of QECC stabilizers to be implemented in constant depth using pairwise XX and ZZ measurements. QECC code conversion schemes [30] can also be implemented locally using pairwise interactions. Additionally, after carefully choosing the pairwise interaction schedules, our method might lead to results similar to the ones from [31]. Our decomposition involves only two-body measurements and may find relevance to subsystem codes such as the Bacon-Shor code [32] or its Floquet version [31].
@@
Conclusion: —
We introduced a novel and very space-time efficient decomposition of Pauli exponentials. Compared to previous methods of implementing long range interactions (e.g. LHZ [14]) our method uses only local two-body interactions, has a linear qubit cost (versus quadratic qubit costs, or time intensive shuttling or SWAP gate schedules) and is applicable to a very wide range of quantum computations, from NISQ to the automatic Floquet-itification of QECC codes. Future work will focus on other fault-tolerance implications of our result.
Acknowledgments: —

### v4

@@ Abstract
We decompose for the first time, under the very restrictive linear nearest-neighbour connectivity, $Z\otimes Z\ldots\otimes Z$ exponentials of arbitrary length into circuits of constant depth using $\mathcal{O}(n)$ ancillae and two-body XX and ZZ interactions. Consequently, a similar method works for arbitrary Pauli exponentials. We prove the correctness of our approach, after introducing novel rewrite rules for circuits which benefit from qubit recycling. The decomposition has a wide variety of applications ranging from the efficient implementation of fault-tolerant lattice surgery computations, to expressing arbitrary stabilizer circuits via two-body interactions only, parallel decoding of quantum error-correcting computations and to reducing the depth of NISQ computations, such as VQE.
I Introduction
@@ I.1 Novelty
Without being generalized to arbitrary rotation angles or number of qubits, constructions similar to ours appeared in [2, 3]. Those constructions were derived and verified by applying the ZX-calculus on small scale diagrams. In contrast, we introduce novel circuit rewrite rules, derive and show the correctness of our decomposition using circuit diagrams. We could have used ZX [18] for illustrating through spider fusions and strong complementarity the correctness of the derivation from Fig. 8, but, because we discovered the FUSE rule by circuit rewrites, it felt more natural to us to stay in the realm of quantum circuits.
@@ II Methods
II.1 Measurement-based implementation
When implementing the Pauli exponentials on fault-tolerant machines, the $\theta$ rotation is first decomposed into a sequence of rotation gates supported by the QEC. This increases the depth proportionally to the length of the gate sequence. Moreover, the depth will additionally increase by a constant factor, as each rotation is probabilistic and might require corrections [4] (e.g. if the X-measurement results in the -1 eigenvalue, then a corrective rotation of $2\theta$ is needed). Corrections can be tracked in a Pauli frame, when the angles correspond to Pauli gates.
@@ II.1 Measurement-based implementation
II.2 Decomposing the Pauli exponentials
@@ III.1 Lattice Surgery
The general compilation strategy for lattice surgery is to first decompose arbitrary algorithms into generalized Pauli exponentials [4], and then to implement these exponentials via multi-body measurements. This approach has multiple disadvantages.
The resulting Pauli exponentials become very long-range and cover large portions of the layout. This hinders the parallelisation of the operations, and alternative compilation methods have been proposed to overcome this, such as [16] (shows how to perform a constant depth CNOT using two-body measurements), and [29], which compiles to Clifford+T instead of exponentiated Paulis.
@@
III.2 Parallel Decoding
Our decomposition enables the parallelisation of decoding by setting an upper bound on the maximum (equivalent) decodable distance that a QEC decoder will need to handle within one QEC cycle. In the following, we introduce the concept of equivalent decodable distance and restrict the discussion, without loss of generality, to rotated surface codes operated by lattice surgery.
An additional disadvantage of long-range Pauli exponentials is the large number of physical qubits (data and syndrome) that need to be decoded. For distance $d$, a surface code patch will require $12d^{2}-$ 1 physical qubits (data and syndrome). The decoding time scales with the size of the surface code patch. Usually, the decoding time is upper bounded by a low-degree polynomial, however, in practice it is still challenging to decode distances larger than 21 within the cycle time of the surface code.
In general, the lattice surgery merge operation between two surface code patches will cover $q=(m+1)(2d^{2}-1)$ qubits, where $m$ is the Manhattan distance between the patches. Decoding the long-range interaction is equivalent to decoding a single patch of distance approximately $\sqrt{\frac{q}{2(m+1)}}$.
As an example, merging two neighboring patches of distance 21 will yield a patch covering $2\times 881=1762$ qubits. This is roughly equivalent to a distance 29 patch. If the patches are separated by a single logical patch of distance 21, the equivalent decoding distance would be 36.
In practice, decoding is very challenging. The computational complexity of the decoding algorithms scales polynomially (more than linear) [30], and surface code compilers do not guarantee short-range implementations of Pauli exponentials. Assuming a logical computation of 100 logical qubits, a compact, resource efficient layout of surface code patches would consist of approximately 150 patches[4]. The average Manhattan distance between two arbitrary patches would be around 25, and the average equivalent decoding distance would be 105. For an instance of Shor’s algorithm with 2048 logical qubits, the best-case equivalent decoding distance would be around 200. For a decoder with cubic runtime complexity [30], this would have an $\approx(10^{2})^{3}=10^{6}$ increase in decoding time compared to decoding distance 21.
Our decomposition sets an upper bound on the equivalent decoding distance of arbitrary Pauli exponentials implemented via lattice surgery. For example, in Fig. 7 the blue interactions always cover only four patches. Assuming the distance of the patches is 21, the maximum equivalent decoding distance to decode would be 42.
At the same time, as seen in Fig. 7, all blue patches can be decoded in parallel. Our parallelisation method is reminiscent of the proposal from [31] (fixing the number of physical qubits per decoder) and the constructions from [32, 33] (decoders which communicate messages). For the latter, the decoders will communicate the observed correlated logical errors [17] after the long-range interaction is split via lattice surgery.
We conjecture that our decomposition, together with the methods from [31, 32, 17] will lead to the same decoding performance for Pauli exponentials of arbitrary length. Future work will focus on this analysis.
III.3 Constant Depth Fault-Tolerant Protocols
Recently, there has been an increased interest in using only two-body interactions to implement error correction (e.g. [3, 2]). Two-body interactions are preferred as quantum computers have limited connectivity between the qubits, and very often the connectivity is not long range (e.g. [19]).
From the perspective of stabilizer QECC codes, our decomposition allows any set of QECC stabilizers to be implemented in constant depth using pairwise XX and ZZ measurements. QECC code conversion schemes [34] can also be implemented locally using pairwise interactions. Additionally, after carefully choosing the pairwise interaction schedules, our method might lead to results similar to the ones from [35]. Our decomposition involves only two-body measurements and may find relevance to subsystem codes such as the Bacon-Shor code [36] or its Floquet version [35].
@@ III.4 NISQ computations
IV Generalization
So far, we considered constant depth decompositions of Pauli strings of the form $Z\otimes...\otimes Z$. Nevertheless, there might be qubits which are not involved in the operation, and the corresponding Pauli string will include an $I$ term for those qubits.
Without loss of generality, we discuss the constant depth, nearest-neighbour implementation of exponentials of the form $Z\otimes\ldots I\ldots\otimes Z$ – a single $Z$ at the start and the end of the string and at least one $I$ in between.
Considering the circuit from Fig. 1, the challenge is to jump over the qubits which should not be operated by $Z$ operators – these qubits are left unchanged by the $I$. In the following, we consider the position (index) of the $Z$ operators in the string (first, last), such that the discussion will revolve around storage and layouts.
One solution is to permute the qubits such that all $I$ terms are moved to the end on the Pauli string. For example, $Z\otimes\ldots I\ldots\otimes Z$ will be implemented via a single two-body nearest-neighbour interaction $Z\otimes Z\ldots\otimes I$. However, sorting introduces two SWAP networks (one before the Pauli string exponential, and another one afterwards) of logarithmic depth (e.g. [9]) instead of a constant depth.
Another solution is to use qudits (moving the qubit state into a higher dimension in order to protect it from qubit $Z$ gates) or special layouts which enable the constant depth construction of Bell pairs using the method from [16].
IV.1 Qudits
It is possible, but not necessarily practical with the current generation of quantum computers, to embed the qubit computation into a system of qudits. To this end, we choose ququarts whose states are spanned by the following basis states: |$\ket{0}$, |$\ket{1}$, |$\ket{2}$, |$\ket{3}$. We will assume that all the wires are ququarts instead of qubits.
We make the observation that qubit gates can be applied on qudits [38] and that whenever this happens, the higher levels of the ququart are left unchanged. For example, the qubit $Z_{2}$ gate (the subscript indicates the dimension of the qudit, e.g. two is for qubits, and four for ququarts) can be embedded into a $4\times 4$ matrix written in block form as $Z_{4}=\big{(}\begin{smallmatrix}Z_{2}&0\\ 0&I_{2}\end{smallmatrix}\big{)}$.
[…]
$I_{2}(a\ket{0}+b\ket{1})=X_{4}^{2}Z_{4}X_{4}^{2}(a\ket{0}+b\ket{1})$
Therefore, whenever ququarts and the $X_{4}$ and $Z_{4}$ gates are supported, we can implement the qubit $Z\otimes\ldots I\ldots\otimes Z$ Pauli string by replacing each $I$ with the single ququart $X_{4}^{2}Z_{4}X_{4}^{2}$. The result is a continuous string of qubit $Z$ operators which can be implemented according to the decomposition from Fig. 1.
An alternative to using ququarts is to use one additional qubit $q$ placed next to each ancilla $a$ from Fig. 1. These $q$ qubits will be initialized in |$\ket{0}$, and the $I$ operators from the Pauli string would be replaced by $SWAP(a,q)$, a $Z$ on $a$, followed by $SWAP(a,q)$. This construction would not be nearest-neighbour, but next-nearest neighbour.
IV.2 The qubit layout
The nearest-neighbour implementation of the two-body interaction is only possible if the qubit layout supports it. For example, in Fig. 7 it is not possible to implement the blue merge operations between nearest-neighbour patches, because each logical Q-patch is surrounded by one ancilla A-patch and three routing patches (not marked by any letter).
In a linear nearest neighbour layout of qubits, the logical qubits are arranged in a line similar to $\begin{smallmatrix}Q&0&Q&0\ldots&Q\end{smallmatrix}$, where 00 are ancillae initialized in |$\ket{0}$. In this scenario, pairwise nearest-neighbour interactions are not possible without resorting to ququarts, as discussed in the previous section. Consequently, the long range construction from [19] uses nearest-neighbour pairwise interactions, but it uses the less general layout $\begin{smallmatrix}Q&0&0&0\ldots&Q\end{smallmatrix}$, where all the qubits between the first and last one are initialized in |$\ket{0}$.
A bi-linear nearest neighbour layout of qubits of the form $\begin{smallmatrix}Q&0&Q&0\ldots&Q\\ r&r&r&r\ldots&r\end{smallmatrix}$ allows for jumping over qubits operated by $I$ and maintaining the nearest-neighbour decomposition – the routing qubits $r$ can be used first to prepare Bell pairs using the constant depth recipe from [16]. The layout from Fig. 7 has the sane properties the bi-linear nearest neighbour layout has.
V Conclusion
We introduced a novel and very space-time efficient decomposition of Pauli exponentials. Compared to previous methods of implementing long range interactions (e.g. LHZ [14]) our method uses only local two-body interactions, has a linear qubit cost (versus quadratic qubit costs, or time intensive shuttling or SWAP gate schedules) and is applicable to a very wide range of quantum computations, from NISQ to the automatic Floquet-itification of QEC codes. Our approach is also enabling the parallelisation and message passing distribution of QEC decoders. Future work will focus on a more detailed analysis of mentioned applications.
VI Acknowledgments